Mike wants to make meatloaf. His recipe uses a total of 7 pounds of meat. If he uses a 3 to 1 ratio, how much pork will he use? Enter in mixed simplest term
Answer by Theo(13342)

since pork is a meat, then i have to assume that he is using 7 pounds of pork.

you are specifying a ratio of 3 to 1, which i am assuming is a meat to other ingredients ratio.

therefore your ratio of pork to other ingredients is assumed to be 3 to 1.

this means you have 3 parts of pork for every 1 part of other ingredients.

if you let p represent the units of pork and o represent the units of other ingredients, then your ratio can be expressed as p/o = 3/1.

you can solve for p to get p = 3 * o.

you can solve for o to get o = 1/3 * p.

since we know what p is, we'll solve for o.

o = 1/3 * p becomes o = 1/3 * 7 which becomes o = 7/3 which becomes o = 2 and 1/3 units of other ingredients.

since the units are pounds, then he needs 2 and 1/3 pounds of other ingredients to mix with 7 pounds of pork.

you can use your calculator to divide 7 by (2 and 1/3) and you will get 3, because 3 * 2 and 1/3 is equal to 6 and 3/3 which is equal to 7.